Jacob decided to ride his bicycle across the country during his 3-month summer vacation. The route he took from Boston to Seattle covered a total of 3760 miles. To keep his mind occupied during some of the long flat stretches of countryside, he often did algebra problems in his head. One day, for example, he determined that his distance from his starting point was exactly 70 miles more than twice the distance remaining until the finishing point. How far was Jacob from the finishing point of his journey when he made that calculation?
Answer by Boreal(15235)

Distance to starting point is x
From there to finishing point is 3760-x
x-70=2(3760-x)
x-70 = 7520-2x
3x=7590
x=2530
He is 2530 miles from his stating point
He is 1230 miles from his finishing point ANSWER
The 2530 miles is 70 more than twice 1230 miles.